In a cribbage game with 2 people, each player is dealt 6 cards to start a round. How many 6- card hands are possible?


1
Expert's answer
2021-06-07T19:31:19-0400

This is a combination of 52 cards taken 6 at a time per person.

"N=52 \\

n=12 \\

52C12 = \frac{52!}{(52-12)!(12!)} \\

= 2.0638 \times 10^{11}"